# Modeling fibrous tissue in vascular fluid-structure interaction: a morphology-based pipeline and biomechanical significance

### v1

@@ Abstract
We propose a suite of technologies for analyzing the interaction between anisotropic arterial walls and blood flow for subject-specific geometries. Utilizing an established lumen modeling strategy, we present a comprehensive pipeline for generating the thick-walled artery models. Through a specialized mesh generation procedure, we obtain the meshes for the arterial lumen and wall with mesh continuity across the interface ensured. Exploiting the centerline information, a series of procedures is introduced for generating local basis vectors within the arterial wall. The procedures are tailored to handle thick-walled and, in particular, aneurysmatic tissues in which the basis vectors may exhibit transmural variations. Additionally, we propose methods to accurately identify the centerline in multi-branched vessels and bifurcating regions. The developed fiber generation method is evaluated against the strategy using linear elastic analysis, demonstrating that the proposed approach yields satisfactory fiber definitions in the considered benchmark. Finally, we examine the impact of anisotropic arterial wall models on the vascular fluid-structure interaction analysis through numerical examples. For comparison purposes, the neo-Hookean model is considered. The first case involves an idealized curved geometry, while the second case studies an image-based abdominal aorta model. The numerical results reveal that the deformation and stress distribution are critically related to the constitutive model of the wall, while the hemodynamic factors are less sensitive to the wall model. This work paves the way for more accurate image-based vascular modeling and enhances the prediction of arterial behavior under physiologically realistic conditions.
@@ 1 Introduction
Over the years, image-based computational fluid dynamics (CFD) and fluid-structure interaction (FSI) gradually get matured and offer a non-invasive approach for quantifying hemodynamic factors, including pressure, wall shear stress (WSS), fractional flow reserve, etc [31, 74, 76]. These factors are crucial for understanding disease progression, making surgical plans, and designing medical devices [73]. As a more physiologically realistic model for vasculature, FSI has been demonstrated to be superior in assessing WSS and related biomarkers [49, 50, 61], characterizing pulse wave propagation [47, 53], and most importantly, analyzing stresses of the tissue wall [12]. The latter plays a critical role in a variety of clinical scenarios, such as risk assessment of aneurysms [64], evaluation the impact of balloon angioplasty on vascular tissue [16], etc.
Despite the encouraging progress in image-based modeling, we witness a dichotomy between the fields of vascular FSI and vascular constitutive modeling. Most existing vascular FSI studies neglect physiological details that are discovered and quantified in biomechanical modeling. For example, isotropic hyperelastic models are still widely used to characterize the mechanical behavior of arterial walls in vascular FSI investigations [11, 15, 16, 49, 50, 54, 69, 67, 79]. In the meantime, the detailed constitutive models that incorporate biomechanics and mechanobiology information have been actively developed [42, 64, 59]. Some recent FSI studies, albeit based on idealized geometries, has revealed the significance of anisotropy, especially in transmural stress distribution [8, 9, 75]. We view the that one major hurdle in the field is the lack of robust tools in characterizing the tissue wall based on physiological facts and image data. This thus becomes the primary research thrust of our work.
@@ 1.1 Physiological background and constitutive modeling
In histology, like most biological tissues, arteries are composed of extracellular matrix (e.g., elastin and collagen), cells (e.g., endothelial and smooth muscle cells), and water. Specifically, the collagen fibers and smooth muscles provide significant stiffness when being stretched and are embedded in groundmatrix which is isotropic and much softer. This composition endowed the tissue with directional and nonlinear attributes. Due to the abundant content of water, most tissues behaves in a nearly-incompressible manner. The arterial wall consists of three layers, characterized by distinct thickness, components, and mechanical properties [26, 37]. For example, the collagen fibers in the media form two helically oriented groups, while the orientation in the other layers are more dispersed.
Informed from physiological knowledge and biomechanical measurements, constitutive models for the artery have been evolved to accurately capture physiological details [21]. Exponential function was introduced to characterize the strain hardening due to collagen fibers [19]. The anisotropy is typically modeled through invariants based on structural tensors [32]. The composite nature of the artery is taken into account by distinguishing the non-collagenous groundmatrix and the collagen fibers, which are described by isotropic and anisotropic hyperelastic models, respectively. One representative work is the one proposed by Holzapfel, Gasser, and Ogden [34], which invokes three parameters for a mechanically homogeneous layer. Through adopting different material parameters, that model can be readily converted to characterize the multi-layer structure of arterial walls. A significant improvement was made through taking the fiber dispersion into account, and it is currently known as the GOH model [26]. Further improvement of that model involves the exclusion of fiber contribution under compression [36] and differentiate the dispersion level in the in-plane and out-of-plane directions [35]. We also mention that the well-possedness of a variety of anisotropic arterial tissue models has been confirmed [10], paving the way towards reliable analysis.
@@ 1.2 Image-based modeling
A mature pipeline exists for preparing necessary image and physiological data for hemodynamic analysis, with the procedures integrated into finite element analysis software [3, 7, 76]. An underlying concept is to construct the vessel wall as two-dimensional (2D) cross-sectional profiles swept along vessel pathlines. Each cross-section profile is generated by 2D image segmentation and is geometrically represented as parameterized closed curves. Through a lofting procedure along the pathline, the closed curves are joined as a NURBS surface that enclose the lumen of interest. Boolean and smooth operations are necessary to join different luminal branches together and remove non-physiological artifacts from the joined surfaces [2, 77]. The eventual geometric model constitute a boundary representation (B-Rep) of the segmented region, which can be discretized by mesh generator for simulation purposes. We mention that the above-described procedure is typical and relatively mature for subject-specific vascular simulation, despite ongoing progress towards three-dimensional (3D) segmentation, that may eventually bypass the above procedure entirely [5, 76].
@@
There are a few critical issues that limit the scope of applicability of the aforementioned modeling strategy. Although the geometric model created is sometimes termed as “solid model”, the word ‘solid’ is a CAD terminology and has nothing to do with the solid tissue. In fact, the most widely adopted imaging modalities, such as computed tomography angiography and magnetic resonance angiography, cannot identify the vascular and surround tissues due to their lack of tissue contrast. The consequent modeled region is the lumen where blood flows. Although people leveraged membrane [24] or shell theories [55] to enable FSI simulations without invoking a 3D arterial tissue model, we need to be aware that the vessel wall thickness can be about 10% of the lumen diameter, rendering the thin-wall assumption questionable. Therefore, generating a vessel wall model is compulsory for vascular FSI. One approach for wall modeling relies on the boundary layer (BL) mesh generation algorithm. By switching the layer growth direction to extrude outward from the vessel wall, the generation of volumetric wall meshes becomes feasible and maintains mesh continuity across the fluid-solid interface [51, 60, 80]. Additionally, by setting the extrusion thickness as a fraction of the local vessel radius, variable wall thickness can be achieved [51]. This method works well for geometrically simple vascular structures, such as those with minimal bifurcations and relatively uniform luminal diameters. However, for arterial trees with complex geometric shapes, self-intersection during the extrusion process becomes unavoidable. In CFD, advanced BL mesh generation algorithms (such as smoothing, contraction, trimming, etc.) may automatically address the self-intersection issues. When using the algorithms to generate the arterial wall geometry, they may result in physiologically unrealistic geometries of the vessel wall.
@@
Also, the anisotropic constitutive model requires the definition of the fiber orientation at points in the tissue. This is by no means a trivial task for realistic arterial models. Several strategies have been proposed in the literature, each with pros and cons. The first is morphology-based and generates a local basis vectors pointed in the circumferential, radial, and axial directions first. The centerline [51, 62] or the tangential planes [45] from the geometric model are utilized for the generation of the local basis vectors. With the knowledge of the angle between the fiber direction and the local basis vectors, the fiber orientations are then explicitly defined. This approach has been applied to investigating carotid bifurcation [45] and abdominal aortic aneurysm [62]. Yet, the considered geometry was relatively simple, and the algorithm needs to be refined to handle multi-branched vessels. In [4], a static linear elastic analysis was performed to extract the principal stress directions, which are assumed to align with the local radial, circumferential, and axial directions. An even simpler method is based on the Laplace equation with prescribed boundary conditions [52]. The local basis vectors for the mean fiber directions are then derived from the gradient of the resulting scalar field. The above approaches rely on the solution of differential equations, which necessitates a proper boundary condition settings.
@@
The second approach assumes that the fibers can realign and reorient in response to external mechanical stimuli with the goal of enhancing its load-bearing capacity. In [29], a biomechanically-driven fiber remodeling method was proposed, in which fiber directions are iteratively updated according to the external load applied. In specific, the fibers are repositioned within a plane spanned by the two maximum principal stress directions, with the angle depending on the ratio of the magnitudes of the two principal stresses. This approach has been applied in studies of carotid bifurcation [30]. In a different approach, a theory was developed to investigate the remodeling of continuously distributed collagen fibers in soft connective tissues, relating the fiber orientation and the mean fiber stretch to the macroscopic stresses within the tissue [23].
@@ 1.3 Contribution and organization of this article
This contribution of this work involves the following aspects. First, we develop an image-based modeling and mesh generation pipeline to generate high-quality subject-specific meshes suitable for FSI analysis. Instead of manipulating the wall surface mesh, we begin by segmenting the exterior wall surface on each slice of 2D images. The exterior surface can then be constructed by lofting, boolean operation, and surface fairing. Existing CAD and polygon mesh processing tools [2, 13] guarantee a flexible and robust way of model editing. With both the interior and exterior surface representation obtained, we may determine the B-Reps of the lumen as well as the vascular wall by closing the planar faces at an inlet and outlets, leading to a discrete mesh suitable for FSI analysis [66, 28]. This pipeline is detailed in Section 2.1, and it allows us to prepare an FSI mesh even when the imaging modality cannot resolve the exterior surface. The introduction of a CAD or a polygon mesh enables editing the wall model to make sure it is smooth and physiologically realistic. This can be especially beneficial when the arterial tree has complex geometries with multiple branches.
Second, we propose a morphology-based approach to define the fiber architecture within the vascular wall in Section 2.2. Our approach relies on the centerline and is refined for complex arterial trees. For thick-walled arterial models, the local basis vectors exhibit transmural variations, and they can be particularly pronounced in aneurysmatic vessels. For multi-branched vessels, multiple centerlines exist, and misidentification of the centerline may occur near bifurcations or when two branches are close to each other. These issues are addressed by a suite of refined strategies for constructing the local basis vectors. Compared to the method that generates local basis vectors using linear elasticity analysis, our approach yields a higher quality result near a bifurcation region, as elaborated in Section 3.1. Based on the local basis vectors, the anisotropic constitutive model for arteries can be implemented for subject-specific geometries, which is discussed in Section 2.3. At the end of Section 2, we present the procedures for generating the prestress within the tissue, which serves as an initialization step for vascular FSI simulations.
Third, we apply the fiber-reinforced anisotropic constitutive model to investigate the vascular FSI problem. We start by considering a simulation of an idealized curved geometry. The preliminary validation results demonstrate good agreement with existing observations. Furthermore, we provide insights into several noteworthy concerns, such as the relationship between fiber angle as well as dispersion in anisotropic arterial models and the FSI results. For specific details, refer to Section 3.2. Subsequently, we perform an FSI analysis of a patient-specific abdominal aortic model, revealing that the WSS distribution of the anisotropic model closely resembles that of the isotropic model. Nevertheless, their stress and strain behaviors qualitatively differ, as discussed in Section 3.3. We draw a conclusion in Section 4.
Figure 1: (a) Creating pathlines along the vessels of interest; (b) creating 2D segmentations along each pathline; (c) constructing a geometric model through lofting the segmentations, and generating the BL mesh based on it; (d)-(f) generating the exterior wall surface utilizing the newly defined contour points; (g) defining the geometric model of the vascular tissue by combining the exterior wall surface with the lumen wall surface.
@@ 2 Methodology
2.1 Imaged-based modeling
In this section, we present a comprehensive image-based modeling approach, which is constructed upon the existing lumen modeling strategy [7, 76]. We begin by reviewing the existing critical steps for lumen modeling, including medical image visualization, extraction of vascular pathlines, 2D segmentation of vessel lumens, construction of vessel boundary surfaces, and mesh generation. Built upon the vessel lumen generation method, we propose a modeling procedure for the arterial wall. Additionally, leveraging the vessel centerline, the generation of the local basis vectors within the multi-branched arterial wall are developed. Finally, we also introduce the generation of prestress in the vascular wall, a necessary initialization procedure for vascular FSI simulations.
@@ 2.1.1 Lumen modeling
The technology for image-based vascular lumen modeling is relatively mature. In the following, we provide a brief overview of the vascular lumen generation procedures that are related to the arterial tissue modeling.
@@
After acquiring medical images, they undergo a preprocessing procedure, involving essential steps of noise reduction, image smoothing, histogram equalization, etc. The objective is to improve the accuracy of subsequent analysis by minimizing artifacts and enhance contrast. After that, 2D slices from the images are stacked to form a volumetric representation, and volume rendering techniques are employed to visualize anatomical structures in three dimensions. This procedure is instrumental in the subsequent vascular path planning.
@@
Segmentation divides medical images into segments (sets of pixels) to locate areas and boundaries of interest. In this study, 2D image segmentation is employed to extract the geometric shapes of blood vessels, providing essential data for subsequent vascular modeling. We identify points close to the center of the vascular lumen to outline the trajectory of the vessels, and those points are termed as path points. Pathlines can then be generated by interpolating the path points, and the planes perpendicular to the pathlines accurately delineate areas apt for segmentation (Figure 1(a)). On each plane, a contour can be generated with extraction methods including level-set [20, 78] or threshold techniques [44, 63] (Figure 1(b)). It is worth noting that if the path points are not chosen sufficiently close to the lumen center, the interpolated vascular pathlines may not accurately represent the true curvature of the vessels, causing it difficult to accurately extract image pixel intensities in the 2D profiles of vessels perpendicular to the path.
@@
    The segmented contour on each plane is typically represented by a series of contiguous line segments. This contour can be refined into a smooth, closed B-spline curve interpolated using a series of contour points, which are sampled from the original contour. Degree elevation and knot insertion are invoked to ensure compatibility of the B-spline curves of different contours.
@@
    To obtain a geometric model that faithfully represents the real blood vessels, it is required to employ smoothing techniques to eliminate non-physiological features that arise near the bifurcation location, particularly when there are sharp variations in the vessel radii of different branches. The widely used smoothing methods include the Laplacian or Taubin methods [13].
The above procedure allows us to obtain physiologically realistic discrete surface of the vascular lumen, and it has been integrated into open-source vascular modeling software [7, 76]. Despite that, the modeling of arterial walls is absent from those software. In the following, we propose a suite of techniques for the modeling of vascular tissue, built upon the lumen surface construction techniques.
Figure 2: (a) Closing the annular planar surfaces to form the vessel wall, followed by the generation of the volumetric mesh for the vessel wall; (b)&(c) combining the meshes from the tissue and lumen to complete the vascular FSI mesh generation.

### v2

@@ Abstract
Modeling fibrous tissue for vascular fluid-structure interaction analysis poses significant challenges due to the lack of effective tools for preparing simulation data from medical images. This limitation hinders the physiologically realistic modeling of vasculature and its use in clinical settings. Leveraging an established lumen modeling strategy, we propose a comprehensive pipeline for generating thick-walled artery models. A specialized mesh generation procedure is developed to ensure mesh continuity across the lumen and wall interface. Exploiting the centerline information, a series of procedures are introduced for generating local basis vectors within the arterial wall. The procedures are tailored to handle thick-walled tissues where basis vectors may exhibit transmural variations. Additionally, we propose methods for accurately identifying the centerline in multi-branched vessels and bifurcating regions. These modeling approaches are algorithmically implementable, rendering them readily integrable into mainstream cardiovascular modeling software. The developed fiber generation method is evaluated against the strategy using linear elastic analysis, demonstrating that the proposed approach yields satisfactory fiber definitions in the considered benchmark. Finally, we examine the impact of anisotropic arterial wall models on the vascular fluid-structure interaction analysis through numerical examples, employing the neo-Hookean model for comparative purposes. The first case involves an idealized curved geometry, while the second studies an image-based abdominal aorta model. Our numerical results reveal that the deformation and stress distribution are critically related to the constitutive model of the wall, whereas hemodynamic factors are less sensitive to the wall model. This work paves the way for more accurate image-based vascular modeling and enhances the prediction of arterial behavior under physiologically realistic conditions.
@@ 1 Introduction
Over the years, image-based computational fluid dynamics (CFD) and fluid-structure interaction (FSI) gradually get matured and offer a non-invasive approach for quantifying hemodynamic factors, including pressure, wall shear stress (WSS), fractional flow reserve, etc [31, 74, 76]. These factors are crucial for understanding disease progression, making surgical plans, and designing medical devices [73]. As a more physiologically realistic model for vasculature, FSI has been demonstrated to be superior in assessing WSS and related biomarkers [49, 50, 61], characterizing pulse wave propagation [47, 53], and most importantly, analyzing stresses of the tissue wall [12]. The latter plays a critical role in a variety of clinical scenarios, such as risk assessment of aneurysms [64], evaluation of the impact of balloon angioplasty on vascular tissue [16], etc.
Despite the encouraging progress in image-based modeling, we witness a dichotomy between the fields of vascular FSI and vascular constitutive modeling. Most existing vascular FSI studies neglect physiological details that are discovered and quantified in biomechanical modeling. For example, isotropic hyperelastic models are still widely used to characterize the mechanical behavior of arterial walls in vascular FSI investigations [11, 15, 16, 49, 50, 54, 67, 69]. In the meantime, the detailed constitutive models that incorporate biomechanics and mechanobiology information have been actively developed [42, 64, 59]. Some recent FSI studies, albeit based on idealized geometries, have revealed the significance of anisotropy, especially in transmural stress distribution [8, 9, 75]. We view the one major hurdle in the field is the lack of robust tools in characterizing the tissue wall based on physiological facts and image data. This thus becomes the primary research thrust of our work.
@@ 1.1 Physiological background and constitutive modeling
In histology, like most biological tissues, arteries are composed of extracellular matrix (e.g., elastin and collagen), cells (e.g., endothelial and smooth muscle cells), and water. Specifically, the collagen fibers and smooth muscles provide significant stiffness when being stretched and are embedded in groundmatrix which is isotropic and much softer. This composition endows the tissue with directional and nonlinear attributes. Due to the abundant content of water, most tissues behave in a nearly incompressible manner. The arterial wall consists of three layers, characterized by distinct thickness, components, and mechanical properties [26, 37]. For example, the collagen fibers in the media form two helically oriented groups, while the orientations in the other layers are more dispersed.
Informed from physiological knowledge and biomechanical measurements, constitutive models for the artery have been evolved to accurately capture physiological details [21]. Exponential function was introduced to characterize the strain hardening due to collagen fibers [19]. The anisotropy is typically modeled through invariants based on structural tensors [32]. The composite nature of the artery is taken into account by distinguishing the non-collagenous groundmatrix and the collagen fibers, which are described by isotropic and anisotropic hyperelastic models, respectively. One representative work was proposed by Holzapfel, Gasser, and Ogden [34], which invokes three parameters for a mechanically homogeneous layer. Through adopting different material parameters, that model can be readily converted to characterize the multi-layer structure of arterial walls. A significant improvement was made by taking the fiber dispersion into account, and it is currently known as the GOH model [26]. Further improvement involves the exclusion of fiber contribution under compression [36] and differentiating the dispersion level in the in-plane and out-of-plane directions [35]. We also mention that the well-possedness of a variety of anisotropic arterial tissue models has been confirmed [10], paving the way towards reliable analysis.
@@ 1.2 Image-based modeling
A mature pipeline exists for preparing the necessary image and physiological data for hemodynamic analysis, with the procedures integrated into finite element analysis software [3, 7, 76]. An underlying concept is to construct the vessel wall as two-dimensional (2D) cross-sectional profiles swept along vessel pathlines. Each cross-section profile is generated by 2D image segmentation and is geometrically represented as closed curves. Through a lofting procedure along the pathline, the curves are joined as a NURBS surface that encloses the lumen of interest. Boolean and smooth operations are necessary to join different luminal branches together and remove non-physiological artifacts from the joined surfaces [2, 77]. The eventual geometric model constitutes a boundary representation (B-Rep) of the segmented region, which can be discretized by a mesh generator for simulation purposes. The above-described procedure is typical and relatively mature for subject-specific vascular simulation, despite ongoing progress towards three-dimensional (3D) segmentation, which may eventually bypass the above procedure entirely [5, 76].
@@
There are a few critical issues that limit the scope of applicability of the aforementioned modeling strategy. Although the geometric model created is sometimes termed as “solid model”, the word “solid” is a CAD terminology and has nothing to do with the solid tissue. In fact, the most widely adopted imaging modalities, such as computed tomography angiography and magnetic resonance angiography, cannot identify the vascular and surrounding tissues due to their lack of tissue contrast. The consequent modeled region is the lumen where blood flows. Although people leveraged membrane [24] or shell theories [55] to enable FSI simulations without invoking a 3D arterial tissue model, we need to be aware that the vessel wall thickness can be about 10% of the lumen diameter [17, 41], rendering the thin-wall assumption questionable. Therefore, generating a vessel wall model is compulsory for vascular FSI. One approach for wall modeling relies on the boundary layer (BL) mesh generation algorithm. By switching the layer growth direction to extrude outward from the vessel wall, the generation of volumetric wall meshes becomes feasible and maintains mesh continuity across the fluid-solid interface [51, 60, 79]. Additionally, by setting the extrusion thickness as a fraction of the local vessel radius, variable wall thickness can be achieved [51]. This method works well for geometrically simple vascular structures, such as those with minimal bifurcations and relatively uniform luminal diameters. However, for arterial trees with complex geometric shapes, self-intersection during the extrusion process becomes unavoidable. Advanced BL mesh generation algorithms (such as smoothing, contraction, trimming, etc.) may automatically address the self-intersection issues. When using the algorithms to generate the arterial wall geometry, they may result in physiologically unrealistic geometries of the vessel wall.
@@
Also, the anisotropic constitutive model requires the definition of the fiber orientation at points in the tissue. This is by no means a trivial task for realistic arterial models. Several strategies have been proposed in the literature, each with pros and cons. The first is morphology-based and generates local basis vectors pointed in the circumferential, radial, and axial directions first. The centerline [51, 62] or the tangential planes [45] from the geometric model are utilized for the generation of local basis vectors. With the knowledge of the angle between the fiber direction and the local basis vectors, the fiber orientations are then explicitly defined. This approach has been applied to investigating carotid bifurcation [45] and abdominal aortic aneurysm [62]. Yet, the considered geometry was relatively simple, and the algorithm needs to be refined to handle multi-branched vessels. In [4], a static linear elastic analysis was performed to extract the principal stress directions, which are assumed to align with the local radial, circumferential, and axial directions. An even simpler method is based on the Laplace equation with prescribed boundary conditions [52]. The local basis vectors for the mean fiber directions are then derived from the gradient of the resulting scalar field. The above approaches rely on the solution of differential equations, which necessitates a proper boundary condition setting.
@@
The second approach assumes that the fibers can realign and reorient in response to external mechanical stimuli with the goal of enhancing theirs load-bearing capacity. In [29], a biomechanically-driven fiber remodeling method was proposed, in which fiber directions are iteratively updated according to the external load applied. In specific, the fibers are repositioned within a plane spanned by the two maximum principal stress directions, with the angle depending on the ratio of the magnitudes of the two principal stresses. This approach has been applied in a study of carotid bifurcation [30]. In a different approach, a theory was developed to investigate the remodeling of continuously distributed collagen fibers in soft connective tissues, relating the fiber orientation and the mean fiber stretch to the macroscopic stresses within the tissue [23].
@@ 1.3 Contribution and organization of this article
This contribution of this work involves the following aspects. First, we develop an image-based modeling and mesh generation pipeline to generate high-quality subject-specific meshes suitable for FSI analysis. Instead of manipulating the wall surface mesh, we exploit CAD and polygon mesh processing tools [2, 13] to deliver a flexible and robust way of model editing. With both the interior and exterior surface representation obtained, we may determine the B-Reps of the lumen as well as the vascular tissue by closing the planar faces at the inlet and outlets, leading to a discrete mesh suitable for FSI analysis [66, 28]. This pipeline is detailed in Section 2.1.
Second, we propose a morphology-based approach to define the fiber architecture within the vascular wall in Section 2.2. Our approach relies on the centerline and is refined for complex arterial trees. Compared to the method based on solving partial differential equations, our approach yields a higher quality result near a bifurcation region, as demonstrated in Section 3.1. Based on the local basis vectors, the anisotropic constitutive model for arteries can be implemented for subject-specific geometries, which is discussed in Section 2.3. At the end of Section 2, we present the procedures for generating the prestress within the tissue, which serves as a necessary initialization step for vascular FSI simulations.
Third, we apply the fiber-reinforced anisotropic constitutive model to investigate the vascular FSI problem. In Section 3.2, we consider a simulation of an idealized curved geometry designed in [8, 9] and analyze the impact of the fiber angle and dispersion on the FSI results. After that, we perform an FSI analysis of a patient-specific abdominal aortic model in Section 3.3, validating the model by comparing with in vivo data and corroborating the observations from the prior idealized geometry. Conclusive remarks are made in Section 4.
@@ 2 Methodology
2.1 Image-based modeling pipeline for vascular lumen and tissue
@@ 2.1.1 Lumen modeling
In the following, we provide a brief overview of the vascular lumen generation procedures [7, 76] that are closely related to the tissue modeling.
@@
After acquiring medical images, they undergo a preprocessing procedure, involving essential steps such as noise reduction, image smoothing, histogram equalization, etc. The objective is to improve the accuracy of subsequent analysis by minimizing artifacts and enhancing contrast. After that, 2D slices from the images are stacked to form a volumetric representation, and volume rendering techniques are employed to visualize anatomical structures in three dimensions.
@@
Segmentation divides medical images into sets of pixels to locate areas and boundaries of interest. In this study, 2D image segmentation is employed to extract the geometric shapes of blood vessels, providing essential data for subsequent vascular modeling. We identify points close to the center of the vascular lumen to outline the trajectory of the vessels, and those points are termed as path points. Pathlines can then be generated by interpolating the path points, and the planes perpendicular to the pathlines accurately delineate areas apt for segmentation (Figure 1 (a)). On each plane, a contour can be generated with extraction methods including level-set [20, 78] or threshold techniques [44, 63] (Figure 1 (b)).
@@
    The segmented contour on each plane is represented by a series of contiguous line segments. This contour can be refined into a smooth, closed B-spline curve interpolated using a series of contour points, which are sampled from the original contour. Degree elevation and knot insertion are invoked to ensure compatibility of the B-spline curves of different contours.
@@
    It is necessary to employ smoothing techniques to eliminate non-physiological features that arise near the bifurcation location, particularly when there are sharp variations in the vessel radii of different branches. The commonly used smoothing methods include the Laplacian or Taubin methods [13].
The above procedure allows one to obtain a physiologically realistic discrete surface of the vascular lumen, and it has been integrated into open-source vascular modeling software [7, 76]. Despite that, the modeling of arterial walls is absent from those software. In the following, we propose a suite of techniques for the modeling of vascular tissue, built upon the lumen surface construction techniques.
Figure 1: (a) Creating pathlines along the vessels of interest; (b) creating 2D segmentations along each pathline; (c) constructing a geometric model and BL mesh; (d) generating the contour points for the exterior surface based on the centroid and the thickness-to-radius ratio $\delta$; (e) creating the closed B-spline curves along each pathline for the tissue exterior surface; (f) constructing the exterior wall surface through lofting; (g) generating the geometric model of the vascular tissue by combining the exterior wall surface with the lumen wall surface.